you are in a airport it takes you one hour longer to flyy to cincinnati at 200mph than it does to return at 250 mph how far away is cincinnati from the airport you are currently in
Answer by mananth(16949)

let cincinati distance be x miles
speed = 200 mph
..
time to cincinnati = x/200
..
return
speed = 250 mph
return time = x/250
..
x/200 - x/250 =1
(250x-200x) / 250*200=1
50x = 250*200
/50
x= 250*200/50
x= 1000 miles
...
m.ananth@hotmail.ca